OpenStudy (anonymous):

Matthew filled his garden with 12 cubic feet of organic dirt. The garden is a rectangular prism with a volume V equals l times w times h where l is length, w is width, and h is depth. If the garden has a depth of 1.5 feet and is 2 feet wide, how long is the garden? 2 feet 3 feet 4 feet 5 feet

OpenStudy (ranga):

V = L * W * H 12 = L * 2 * 1.5 Solve for L